Direct Labor Variances Greeson Clothes Company produced 15,000 units during June of the current year. The Cutting Department used 2,900 direct labor hours at an actual rate of $12.10 per hour. The Sewing Department used 4,800 direct labor hours at an actual rate of $11.80 per hour. Assume that there were no work in process inventories in either department at the beginning or end of the month. The standard labor rate is $12.00. The standard labor time for the Cutting and Sewing departments is 0.20 hour and 0.30 hour per unit, respectively. a. Determine the direct labor rate, direct labor time, and total direct labor cost variance for the Cutting Department and Sewing Department. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.
Answer #1

Calculate following

Cutting Sewing
Labor rate variance (12-12.10)*2900 = 290 U (12-11.80)*4800 = -960 F
Labor efficiency variance (3000-2900)*12 = -1200 F (4500-4800)*12 = 3600 U
Total direct labor cost variance 910 -F 2640 U